# Skyrmion motion in a synthetic antiferromagnet driven by asymmetric spin wave emission

### v1

@@ Abstract
Skyrmions have been proposed as new information carriers in racetrack memory devices. To realise such devices, a small size; high speed of propagation; and minimal skyrmion Hall angle are required. Synthetic antiferromagnets (SAFs) present the ideal materials system to realise these aims. In this work, we use micromagnetic simulations to propose a new method for manipulating them using exclusively global magnetic fields. An out-of-plane microwave field induces oscillations in the skyrmions radius which in turn emits spin waves. When a static in-plane field is added, this breaks the symmetry of the skyrmions and causes asymmetric spin wave emission. This in turn drives motion of the skyrmions, with the fastest velocities observed at the frequency of the intrinsic out-of-phase breathing mode of the pair of skyrmions. This behaviour is investigated over a range of experimentally realistic antiferromagnetic interlayer exchange coupling strengths, and the results compared to previous works studying similar motion driven with an oscillating electric field. Through this the true effect of varying the exchange coupling strength is determined, and greater insight is gained into the mechanism of skyrmion motion. These results will help to inform the design of future novel computing architectures based on the dynamics of skyrmions in synthetic antiferromagnets.
@@ I Introduction
Synthetic antiferromagnets—where two individually ferromagnetic layers are sandwiched around a nonmagnetic spacer layer and coupled antiferromagnetically [22]—have been predicted to be an advantageous platform for skyrmions [23, 21]. This is because of their proposed smaller size, faster velocity, and increased resistance to stray fields compared to skyrmions in ferromagnetic systems [24]. Skyrmions with radii as small as $\sim$50 nm have been experimentally in synthetic antiferromagnets [25, 26]. Skyrmions and domain walls have been shown to move at substantially faster velocities than their ferromagnetic equivalents of up to 900 m/s [27] in the case of skyrmions, or 750 m/s in the case of domain walls [28]. Domain walls have also been shown to depin at current densities as low as $\mathrm{1\times 10^{11}\ A/m^{2}}$ [29, 30]. Due to the challenges in detection, much of the body of research on skyrmions in synthetic antiferromagnets remains theoretical [31], and experimental measurements of skyrmion motion are extremely limited [26, 27, 32].
@@
Skyrmions have been shown to exhibit microwave resonances when excited by magnetic and electric fields [33]. This was originally predicted based on numerical modelling of skyrmions in bulk systems [34], and was first demonstrated experimentally in Cu2OSeO3 [35]. The breathing modes of skyrmions were first simulated in thin films by Kim et al. [36], and have recently been demonstrated experimentally in ferromagnetic thin films [37, 38]. In synthetic antiferromagnets, micromagnetic simulations have been used to demonstrate both the breathing modes [39, 40] and corresponding rotational modes [41]. By exciting these breathing modes using either electric, or magnetic fields, it has been demonstrated that skyrmions in ferromagnetic systems will emit spin waves asymmetrically in a static in-plane field, and this stimulates skyrmion motion with a peak in velocity as a function of frequency that is coupled to the peak of the breathing mode amplitude [16, 42]. This has also been demonstrated in synthetic antiferromagnets where microwave electric fields were used to excite straight-line skyrmion motion [43]. In that work, a dependence of the velocity on the strength of the antiferromagnetic interlayer exchange coupling was demonstrated, however the link between this and the skyrmion breathing modes was not discussed [43]. This microwave field driven approach to skyrmion motion has the advantage that—unlike the conventional method of using an electric current—it does not require a conductive sample, and so can be used to manipulate skyrmions in insulating materials.
@@
In this work, we use microwave magnetic fields in micromagnetic simulations to drive skyrmion breathing modes in a synthetic antiferromagnet, and combine these with an in-plane, static field to excite skyrmion motion in a wiggle-like fashion. We investigate this motion as a function of frequency, and relate it to the intrinsic breathing modes of these skyrmions. We use this understanding of the coupling between velocity and breathing mode frequency to understand the true relationship between the skyrmion velocity and the antiferromagnetic interlayer exchange coupling strength. Finally, we use the additional knowledge of this mechanism to discuss the differences between the previous electric-field driven skyrmion motion [43] and our work. The additional understanding of the relationship between the skyrmion breathing mode and the velocity can be used in applications to modulate the speed of the skyrmion using only the microwave frequency. The emission of spin waves by a domain wall driven by microwave magnetic fields has also very recently been demonstrated [44]. This, together with the similarly recent imaging of spin waves in a synthetic antiferromagnet [45], paves the way to the experimental realisation of our method.
@@ II Methods
The simulation is set up with exchange stiffness $A=15\times 10^{-12}$ J/m, saturation magnetisation $M_{\mathrm{s}}=580\times 10^{3}$ A/m, uniaxial anisotropy constant $K=7\times 10^{5}\ \mathrm{J/m^{3}}$, interfacial DMI $D=3\times 10^{-3}\ \mathrm{J/m^{2}}$ and Gilbert damping parameter $\alpha=0.01$. These values represent typical generic parameters of metallic multilayers with DMI that are usually chosen for synthetic antiferromagnets hosting skyrmions [47]. They are also chosen specifically to mirror those used in ref. [43] in order to better compare results. The system was initialised with an in plane x-y cellsize of 1.5625 nm and an out-of-plane z cellsize of 0.4 nm. The configuration of cells was $256\times 256\times 4$ cells, for a total layer width and length of 400 nm and a total stack thickness of 1.6 nm. Only two of the layers were magnetic, with the remaining two layers not contributing to the simulation other than to mirror the spatial separation of the magnetic layers as in a real system.
@@
To excite the skyrmion motion—unless otherwise specified—a static in-plane field of 0.5 T is used, together with a static out-of-plane field of 50 mT and an out-of-plane microwave excitation field of 20 mT. The microwave excitation field is smaller than that used by other works [15] but a comparable order of magnitude. In order to track the motion of the skyrmion, for each timestep we calculate the guiding centre $\mathbf{R}_{\mathrm{Sk}}$ of each magnetic layer, given by
@@ III.1 Effect of in-plane field on skyrmions
The effect of a static in-plane field on the skyrmion radius is shown in Fig. 1(b-e) for an antiferromagnetic interlayer exchange coupling strength of $-3\times 10^{-4}\ \mathrm{J/m^{2}}$ in a static out-of-plane field of 50 mT. For this antiferromagnetic interlayer exchange coupling strength, the interlayer coupling is sufficiently strong to overcome the effects of the out-of-plane field and so the two skyrmions have equal radius in zero in-plane field $H_{\mathrm{IP}}$. As the field increases above $\sim$0.3 T, the skyrmions in each layer begin to both stretch and shrink. This shrinking is caused by the in-plane field pulling the cores of the two skyrmions apart laterally—as shown in Fig. 1(b,c)—which introduces a breaking of the antiferromagnetic interlayer coupling between spins inside the skyrmion. Thus, the skyrmion size shrinks to minimize this. Because of this, the stretching of the skyrmion is better demonstrated by the ratio between $r_{x}$ and $r_{y}$, as shown in Fig. 1(f). Above fields of $\sim$0.3 T this behaviour starts to change, and steadily changes from 0.5 T onwards, indicating the stretching of the two skyrmions along the direction of the field. On this basis, unless otherwise specified, an in-plane field of 0.5 T is chosen for further simulations which causes a modest deformation of the skyrmion, and also allows for better intercomparison of results with ref. [43]. Simulations of the full hysteresis loops are shown in supplementary note LABEL:supp:note:hyst_loops. The saturation field along the x- and y-axes is 4.3 T, demonstrating that this in-plane field is only a modest $\sim$10 % of saturation.
@@ IV.1 Skyrmion motion driven by the addition of an in-plane magnetic field
Figure 3: In-plane field-driven skyrmion motion in a field of 0.5 T along x, driven at a frequency of 57.4 GHz. Panel (a) shows the x- and y-positions of the skyrmion guiding centre for both the skyrmions in layers 1 and 2 of the SAF. Panel (b) shows the calculated velocities of these guiding centre x- and y-positions as a function of time. Points on the lines in panels (a) and (b) are deliberately undersampled to see both the datapoints of layers 1 and 2. Panel (c) shows the trajectory of both skyrmions, with the y-positions of each skyrmion plotted against the x-position. The colour gradient of each line denotes the time, which is shown on the colour bar to the right of the panel. Arrow shows the direction of the in plane static field.
Considering the static and dynamic characterisation of skyrmions in the simulated system described in Sec. III.1 and Sec. III.2, we now move on to describing the effect of an in-plane field together with an excitation field on the motion of the SAF skyrmion. The frequencies of the two peaks of Fig. 2(a) are found by fitting the data to a Lorentzian lineshape to be 14.3 GHz and 57.4 GHz. A sinusoidal driving field is then used at the higher frequency to drive the skyrmion together with a static out-of-plane field of 50 mT as discussed in Sec. III.2. This is also done with an in-plane static field of 0.5 T. The results of this are shown in Fig. 3. Fig. 3(a) shows the x- and y-positions of the guiding centres of the skyrmions in each layer as a function of time for 20 ns. We see that—while the skyrmion moves steadily along the y-axis at a seemingly constant rate—the x-axis motion proceeds in an oscillatory fashion with an oscillation frequency of $\sim$1 GHz, with an additional very small net movement along the x-axis. This can also be visualised with the trajectories of the two skyrmions shown in Fig. 3(c). We see that the skyrmions move in a wiggle-like fashion, the first report of this kind of motion in a synthetic antiferromagnet. This is also unlike the motion observed in ref. [43] but similar to the trajectories in similar ferromagnetic systems reported in [42]. The position as a function of time is used to calculate the x- and y-velocities of each skyrmion as a function of time, which is shown in Fig. 3(b). The y velocities ramp up from 0 for the first $\sim$2 ns, before maintaining a steady value of $\sim$6 m/s for the rest of the simulation time. Meanwhile, the x velocities oscillate around 0, driving the wiggle-like motion. To calculate the average velocities presented later, an average is taken of the y-axis velocity once it reaches its steady value, here around 3 ns into the simulation time.
@@ IV.2 Skyrmion velocity coupled to out-of-phase breathing
Figure 4: Velocities of the two skyrmions as a function of frequency of the driving field. Datapoints are deliberately undersampled so that both datasets are visible. The intensity of the breathing modes from Fig. 2 is overlaid with grey shading and rescaled to match the peak amplitudes. Inset shows the same schematic from Fig. 1 to illustrate the directions of the magnetic fields and the direction of skyrmion motion.
@@
Between frequencies of $\sim$70 and $\sim$80 GHz in Fig. 4(a)—above the out-of-phase breathing resonant frequency—there is a second broader peak in velocity that is seemingly not related to any skyrmion breathing mode. This is reminiscent of the higher frequency hybridisation of skyrmion breathing with spin wave modes of ferromagnetic [36] and synthetic antiferromagnetic [39] nanodots that has been previously observed. However, as discussed in Sec. II, in this simulation periodic boundary conditions are used. This excludes such spin wave resonant modes, since they originate from spin canting at the edges of the simulation, which the application of periodic boundary conditions removes. We instead suggest that this higher frequency broad peak in velocity arises from exciting a hybrid mode between the skyrmion breathing and a spin wave mode arising from the canting of spins in the applied in-plane field. Because of the motion that would be induced by performing the simulations used to generate Fig. 2 in an in-plane field, this is difficult to show directly. However, in supplementary note LABEL:supp:sec:high_freq_mode we show and discuss simulation of the PSD in the presence of an in-plane field but without a skyrmion, which yields a peak of similar centre frequency and width to this second peak in velocity.
@@ IV.3 The effect of the spin wave emission amplitude
It has been discussed that the motion of the skyrmions in these systems is driven by spin wave emission [42, 16, 43], however, the mechanism of this spin wave emission has not been investigated in detail. This emission is represented more clearly in Figs. 5(a,b) which show the skyrmion when driven at 57.4 GHz in a 0.5 T in-plane magnetic field at 0.2 ns after the start of the simulation; at the onset of skyrmion motion. There is some interference present because of the periodic boundary conditions. Similarly Figs. 5(g,h) show the same situation when the system is driven with a field of 14.3 GHz, the frequency of the in-phase breathing mode. These images are complemented by supplementary videos S1 and S2 which show this spin wave emission over the first 1 ns of the simulation in the presence of an in-plane magnetic field. To further illustrate the spin-wave emission, each of these pairs of colormaps are complemented with line cuts that are take through the centre of the skyrmion, shown in Fig. 5(c,d) for the 57.4 GHz case, and Fig. 5(i,j) for the 14.3 GHz case. These line cuts are in turn complemented by supplementary videos S3 and S4, which again show this profile over the first 1 ns of the simulation, in an in-plane magnetic field of 0.5 mT, driven by an out-of-plane oscillating field of 20 mT, at 57.4 GHz and 14.3 GHz in supplementary videos S3 and S4, respectively. These videos show that the spin waves are emitted at the maxima of the oscillations in the skyrmion radius. This is remarkably similar to the experimental demonstration of spin-wave emission from an oscillating domain wall [44], where the oscillating wall ‘knocks’ the domain interior, stimulating spin waves [44]. It is clear from Fig.5(a,b) and Fig. 5(g,h) that the main difference between being driven at 57.4 GHz—where there is skyrmion motion—and at 14.3 GHz—where there is no skyrmion motion—is the amplitude of the emitted spin waves. It is well understood that the oscillations in radius at the out-of-phase breathing mode are considerably larger than at the in-phase breathing mode [39, 40]. We suggest that the primary factor determining the strength of the emitted spin waves is the magnitude of the change in radius. This is in turn affected by the magnitude of the driving field. However, for the same magnetic field amplitude, this will increase when tuning onto a skyrmion breathing mode resonant frequency, and will also be considerably larger for the out-phase breathing mode frequency.
@@
To investigate this in more detail, we examine the amplitude of the variation in radius in the absence of an in-plane magnetic field. We do this because the deformation of the skyrmions—as shown in Fig. 1—will affect being able to accurately estimate the change in radius. Fig. 5(e,f) show snapshots of the same slice through the centre of the skyrmion at the same time as Fig. 5(c,d) at the same driving field frequency of 57.4 GHz, but without the in-plane field. Likewise, Fig. 5(k,l) show the same, but at a driving field frequency of 14.3 GHz, equivalent to Fig. 5(i,j) but without the in-plane magnetic field. Both of these sets of snapshots are supported by supplementary videos S5 and S6 which show the same data over the first 1 ns of the simulation. For a detailed description of all of the supplementary videos see supplementary note LABEL:supp:sec:videos.
@@
In Fig. 5(m,n), we show a simultaneous study of the skyrmion velocity (Fig. 5(m)) and the magnitude of the variation in radius (Fig. 5(n)), both as a function of the magnitude of the driving microwave magnetic field at a frequency of 57.4 GHz. For details of how the average change in radius was calculated, see supplementary note LABEL:supp;sec:rad_var. In Fig. 5(m), at fields above $\sim$20 mT there is a clear linear relationship between the velocity and the driving field. However, below fields of $\sim$20 mT, the skyrmion velocity drops off faster towards 0 m/s, and there is no motion below driving fields of $\sim$3 mT. This region of no skyrmion motion is represented with the grey shaded box. Since this is a simulated system with no pinning, this is clearly an intrinsic effect of their being an insufficient spin-wave emission amplitude to drive motion. Likewise, examining the magnitude of the variation in radius in Fig. 5(n), we draw the same threshold field for skyrmion motion from Fig 5(m) also as a grey shaded box. For these magnetic parameters, we are then able to define a variation in radius of the skyrmions below which there is no motion of $\langle\Delta r_{\mathrm{Sk}}\rangle$ = 0.44 nm, or $\sim$3.5 % of the static skyrmion radius $r_{0}$.

### v2

@@ Abstract
Skyrmions have been proposed as new information carriers in racetrack memory devices. To realise such devices, a small size; high speed of propagation; and minimal skyrmion Hall angle are required. Synthetic antiferromagnets (SAFs) present the ideal materials system to realise these aims. In this work, we use micromagnetic simulations to propose a new method for manipulating them using exclusively global magnetic fields. An out-of-plane microwave field induces oscillations in the skyrmions radius which in turn emits spin waves. When a static in-plane field is added, this breaks the symmetry of the skyrmions and causes asymmetric spin wave emission. This in turn drives motion of the skyrmions, with the fastest velocities observed at the frequency of the intrinsic out-of-phase breathing mode of the pair of skyrmions. This behaviour is investigated over a range of experimentally realistic antiferromagnetic interlayer exchange coupling strengths, and the results compared to previous works. Through this we demonstrate the true effect of varying the exchange coupling strength, and gain greater insight into the mechanism of skyrmion motion. In an attempt to more accurately reproduce experimental SAF samples, we investigate the effect of changing the ratio between the layer magnetisations in the SAF, and demonstrate that maximum velocity can be achieved when there is 100 % compensation. We also investigate the differences between excitation mechanism, using both electric fields and spin transfer torques to excite skyrmion motion. These results will help to inform the design of future novel computing architectures based on the dynamics of skyrmions in synthetic antiferromagnets.
@@ I Introduction
Synthetic antiferromagnets—where two individually ferromagnetic layers are sandwiched around a nonmagnetic spacer layer and coupled antiferromagnetically [22]—have been predicted to be an advantageous platform for skyrmions [23, 21]. This is because of their proposed smaller size, faster velocity, and increased resistance to stray fields compared to skyrmions observed in ferromagnetic systems [24]. Skyrmions with radii as small as $\sim$50 nm have been experimentally in synthetic antiferromagnets [25, 26]. Skyrmions and domain walls have been shown to move at substantially faster velocities than their ferromagnetic equivalents of up to 900 m/s [27] in the case of skyrmions, or 750 m/s in the case of domain walls [28]. Domain walls have also been shown to depin at current densities as low as $\mathrm{1\times 10^{11}\ A/m^{2}}$ [29, 30]. Due to the challenges in detection, much of the body of research on skyrmions in synthetic antiferromagnets remains theoretical [31], and experimental measurements of skyrmion motion are extremely limited [26, 27, 32].
@@
Skyrmions have been shown to exhibit microwave resonances when excited by magnetic and electric fields [33]. This was originally predicted based on numerical modelling of skyrmions in bulk systems [34], and was first demonstrated experimentally in Cu2OSeO3 [35]. The breathing modes of skyrmions were first simulated in thin films by Kim et al. [36], and have recently been demonstrated experimentally in ferromagnetic thin films [37, 38]. In synthetic antiferromagnets, micromagnetic simulations have been used to demonstrate both the breathing modes [39, 40] and corresponding rotational modes [41]. The breathing and rotational modes of skyrmions have generated interest for magnonics, where, by exciting at the resonant frequency of these modes, they have been used as spin-wave emitters [42, 43, 44]. Building on this, by exciting these breathing modes in the presence of an in-plane magnetic field, it has been demonstrated that skyrmions in ferromagnetic systems will emit spin waves asymmetrically. This stimulates skyrmion motion with a peak in velocity as a function of frequency that is coupled to the breathing mode resonant frequency [16, 45]. In the presence of the in-plane field, there will be a net momentum transfer from the skyrmion to the spin waves, which will cause the skyrmion to move through the conservation of linear momentum [45, 48]. This has also been demonstrated in synthetic antiferromagnets where microwave electric fields were used to excite straight-line skyrmion motion [46]. In that work, a dependence of the velocity on the strength of the antiferromagnetic interlayer exchange coupling was demonstrated, however the link between this and the skyrmion breathing modes was not discussed [46]. This microwave field driven approach to skyrmion motion has the advantage that—unlike the conventional method of using an electric current—it does not require a conductive sample, and so can be used to manipulate skyrmions in insulating materials.
@@
In this work, we use microwave magnetic fields in micromagnetic simulations to drive skyrmion breathing modes in a synthetic antiferromagnet, and combine these with an in-plane, static field to excite skyrmion motion in a wiggle-like fashion. We investigate this motion as a function of frequency, and relate it to the intrinsic breathing modes of these skyrmions. We use this understanding of the coupling between velocity and breathing mode frequency to understand the true relationship between the skyrmion velocity and the antiferromagnetic interlayer exchange coupling strength. Finally, we use the additional knowledge of this mechanism to discuss the differences between the previous electric-field driven skyrmion motion [46] and our work. The additional understanding of the relationship between the skyrmion breathing modes and the velocity can be used in applications to modulate the speed of the skyrmion using only the microwave frequency, and has particular relevance to the emerging field of skyrmion magnonics [47]. The emission of spin waves by a domain wall driven by microwave magnetic fields has also very recently been demonstrated [48]. This, together with the similarly recent imaging of spin waves in a synthetic antiferromagnet [49], paves the way to the experimental realisation of our method.
@@ II Methods
The simulation is set up with exchange stiffness $A=15\times 10^{-12}$ J/m, saturation magnetisation $M_{\mathrm{s}}=580\times 10^{3}$ A/m, uniaxial anisotropy constant $K=7\times 10^{5}\ \mathrm{J/m^{3}}$, interfacial DMI $D=3\times 10^{-3}\ \mathrm{J/m^{2}}$ and Gilbert damping parameter $\alpha=0.01$. These values represent typical generic parameters of metallic multilayers with DMI that are usually chosen for synthetic antiferromagnets hosting skyrmions [51]. They are also chosen specifically to mirror those used in ref. [46] in order to better compare results. The system was initialised with an in plane x-y cellsize of 1.5625 nm and an out-of-plane z cellsize of 0.4 nm. The configuration of cells was 256×256×4256$\times$ 256$\times$ 4256 × 256 × 4 cells, for a total layer width and length of 400 nm and a total stack thickness of 1.6 nm. Only two of the layers were magnetic, with the remaining two layers not contributing to the simulation other than to mirror the spatial separation of the magnetic layers as in a real system.
@@
To excite the skyrmion motion—unless otherwise specified—a static in-plane field of 0.5 T is used, together with a static out-of-plane field of 50 mT and an out-of-plane microwave excitation field of 20 mT. The static in-plane field is chosen to be 0.5 T to be consistent with previous work studying skyrmion motion driven by excitation with an electric field [46]. Likewise the choice of the 50 mT static out-of-plane field is in line with previous work on skyrmion breathing driven by a magnetic field [39, 40]. As discussed in Ref. [39], this small static field serves to aid in the observation of the in-phase skyrmion breathing mode, but does not otherwise alter the behaviour of the breathing modes for fields much smaller than the saturation field, as is the case for our chosen value (see Supplementary Note S1 for measurement of the saturation field.). The microwave excitation field is smaller than that used by other works [15] but a comparable order of magnitude. In order to track the motion of the skyrmion, for each timestep we calculate the guiding centre $\mathbf{R}_{\mathrm{Sk}}$ of each magnetic layer, given by
@@ III.1 Effect of in-plane field on skyrmions
The effect of a static in-plane field on the skyrmion radius is shown in Fig. 1(b-e) for an antiferromagnetic interlayer exchange coupling strength of $-3\times 10^{-4}\ \mathrm{J/m^{2}}$ in a static out-of-plane field of 50 mT. For this antiferromagnetic interlayer exchange coupling strength, the interlayer coupling is sufficiently strong to overcome the effects of the out-of-plane field and so the two skyrmions have equal radius in zero in-plane field $H_{\mathrm{IP}}$. As the field increases above $\sim$0.3 T, the skyrmions in each layer begin to both stretch and shrink. This shrinking is caused by the in-plane field pulling the cores of the two skyrmions apart laterally—as shown in Fig. 1(b,c)—which introduces a breaking of the antiferromagnetic interlayer coupling between spins inside the skyrmion. Thus, the skyrmion size shrinks to minimize this. Because of this, the stretching of the skyrmion is better demonstrated by the ratio between $r_{x}$ and $r_{y}$, as shown in Fig. 1(f). Above fields of $\sim$0.3 T this behaviour starts to change, and steadily changes from 0.5 T onwards, indicating the stretching of the two skyrmions along the direction of the field. On this basis, unless otherwise specified, an in-plane field of 0.5 T is chosen for further simulations which causes a modest deformation of the skyrmion, and also allows for better intercomparison of results with Ref. [46]. Simulations of the full hysteresis loops are shown in supplementary note S1. The saturation field along the x- and y-axes is 4.3 T, demonstrating that this in-plane field is only a modest $\sim$10 % of saturation.
@@ IV.1 Skyrmion motion driven by the addition of an in-plane magnetic field
Figure 3: In-plane field-driven skyrmion motion in a field of 0.5 T along x, driven at a frequency of 57.4 GHz. Panel (a) shows the x- and y-positions of the skyrmion guiding centre for both the skyrmions in layers 1 and 2 of the SAF. Panel (b) shows the calculated velocities of these guiding centre x- and y-positions as a function of time. Points on the lines in panels (a) and (b) are deliberately undersampled to see both the datapoints of layers 1 and 2. Panel (c) shows the trajectory of both skyrmions, with the y-positions of each skyrmion plotted against the x-position. The colour gradient of each line denotes the time, which is shown on the colour bar to the right of the panel. The two insets show the position of the skyrmion in one of the layers at the start (t = 0 ns) and end (t = 10 ns) of the simulation. Scale bar is 100 nm. The arrow in the bottom right shows the direction of the in plane static field.
Considering the static and dynamic characterisation of skyrmions in the simulated system described in Sec. III.1 and Sec. III.2, we now move on to describing the effect of an in-plane field together with an excitation field on the motion of the SAF skyrmion. The frequencies of the two peaks of Fig. 2(a) are found by fitting the data to a Lorentzian lineshape to be 14.3 GHz and 57.4 GHz. A sinusoidal driving field is then used at the higher frequency to drive the skyrmion together with a static out-of-plane field of 50 mT as discussed in Sec. III.2. This is also done with an in-plane static field of 0.5 T. The results of this are shown in Fig. 3. Fig. 3(a) shows the x- and y-positions of the guiding centres of the skyrmions in each layer as a function of time for 20 ns. We see that—while the skyrmion moves steadily along the y-axis at a seemingly constant rate—the x-axis motion proceeds in an oscillatory fashion with an oscillation frequency of $\sim$1 GHz, with an additional very small net movement along the x-axis. This can also be visualised with the trajectories of the two skyrmions shown in Fig. 3(c), together with the insets showing the initial and final positions of the skyrmion. We see that the skyrmions move in a wiggle-like fashion, the first report of this kind of motion in a synthetic antiferromagnet. This is also unlike the motion observed in ref. [46] but similar to the trajectories in similar ferromagnetic systems reported in [45]. The position as a function of time is used to calculate the x- and y-velocities of each skyrmion as a function of time, which is shown in Fig. 3(b). The y velocities ramp up from 0 for the first $\sim$2 ns, before maintaining a steady value of $\sim$6 m/s for the rest of the simulation time. Meanwhile, the x velocities oscillate around 0, driving the wiggle-like motion. To calculate the average velocities presented later, an average is taken of the y-axis velocity once it reaches its steady value, here around 3 ns into the simulation time.
@@ IV.2 Skyrmion velocity coupled to out-of-phase breathing
Figure 4: Velocities of the two skyrmions as a function of frequency of the driving field. Datapoints are deliberately undersampled so that both datasets are visible. The intensity of the breathing modes from Fig. 2 is overlaid with light grey shading and rescaled to match the peak amplitudes. The power spectral density of the spin wave mode arising from the canted spins in the IP field (See supplementary note S2) is also overlaid in a dark grey shading. Inset shows the same schematic from Fig. 1 to illustrate the directions of the magnetic fields and the direction of skyrmion motion.
@@
Between frequencies of $\sim$70 and $\sim$80 GHz in Fig. 4(a)—above the out-of-phase breathing resonant frequency—there is a second broader peak in velocity that is seemingly not related to any skyrmion breathing mode. This is reminiscent of the higher frequency hybridisation of skyrmion breathing with spin wave modes of ferromagnetic [36] and synthetic antiferromagnetic [39] nanodots that has been previously observed. However, as discussed in Sec. II, in this simulation periodic boundary conditions in the x-y plane are used. This excludes such spin wave resonant modes, since they originate from spin canting at the edges of the simulation, which the application of periodic boundary conditions removes. We instead suggest that this higher frequency broad peak in velocity arises from exciting a hybrid mode between the skyrmion breathing and a spin wave mode arising from the canting of spins in the applied in-plane field. Because of the motion that would be induced by performing the simulations used to generate Fig. 2 in an in-plane field, this is difficult to show directly. However, in supplementary note S2 we show and discuss simulation of the PSD in the presence of an in-plane field but without a skyrmion, which yields a peak of similar centre frequency and width to this second peak in velocity. The power spectral density arising from this is shown in Figure 4 as the darker gray shaded region, with a peak at higher frequencies, approximately matching the corresponding peak in velocity. This is discussed in detail in supplementary note S2, however it clearly demonstrates that the peak in velocity arises from hybridisation between the skyrmion breathing mode and the spin-wave resonance of the canted spins when an in-plane field is introduced.
@@ IV.3 The effect of the spin wave emission amplitude
It has been discussed that the motion of the skyrmions in these systems is driven by spin wave emission [45, 16, 46], however, the mechanism of this spin wave emission has not been investigated in detail. This emission is represented more clearly in Figs. 5(a,b) which show the skyrmion when driven at 57.4 GHz in a 0.5 T in-plane magnetic field at 0.2 ns after the start of the simulation; at the onset of skyrmion motion. There is some interference present because of the periodic boundary conditions. Similarly Figs. 5(g,h) show the same situation when the system is driven with a field of 14.3 GHz, the frequency of the in-phase breathing mode. These images are complemented by supplementary videos S1 and S2 which show this spin wave emission over the first 1 ns of the simulation in the presence of an in-plane magnetic field. To further illustrate the spin-wave emission, each of these pairs of colormaps are complemented with line cuts that are take through the centre of the skyrmion, shown in Fig. 5(c,d) for the 57.4 GHz case, and Fig. 5(i,j) for the 14.3 GHz case. These line cuts are in turn complemented by supplementary videos S3 and S4, which again show this profile over the first 1 ns of the simulation, in an in-plane magnetic field of 0.5 T, driven by an out-of-plane oscillating field of 20 mT, at 57.4 GHz and 14.3 GHz in supplementary videos S3 and S4, respectively. These videos show that the spin waves are emitted at the maxima of the oscillations in the skyrmion radius. This is remarkably similar to the experimental demonstration of spin-wave emission from an oscillating domain wall [48], where the oscillating wall ‘knocks’ the domain interior, stimulating spin waves [48]. It is clear from Fig.5(a,b) and Fig. 5(g,h) that the main difference between being driven at 57.4 GHz—where there is skyrmion motion—and at 14.3 GHz—where there is no skyrmion motion—is the amplitude of the emitted spin waves. It is well understood that the oscillations in radius at the out-of-phase breathing mode are considerably larger than at the in-phase breathing mode [39, 40]. We suggest that the primary factor determining the strength of the emitted spin waves is the magnitude of the change in radius. This is in turn affected by the magnitude of the driving field. However, for the same magnetic field amplitude, this will increase when tuning onto a skyrmion breathing mode resonant frequency, and will also be considerably larger for the out-phase breathing mode frequency.
@@
To investigate this in more detail, we examine the amplitude of the variation in radius in the absence of an in-plane magnetic field. We do this because the deformation of the skyrmions—as shown in Fig. 1—will affect the accuracy of determining the change in radius. Fig. 5(e,f) show snapshots of the same slice through the centre of the skyrmion at the same time as Fig. 5(c,d) at the same driving field frequency of 57.4 GHz, but without the in-plane field. Likewise, Fig. 5(k,l) show the same, but at a driving field frequency of 14.3 GHz, equivalent to Fig. 5(i,j) but without the in-plane magnetic field. Both of these sets of snapshots are supported by supplementary videos S5 and S6 which show the same data over the first 1 ns of the simulation. For a detailed description of all of the supplementary videos see supplementary note S3.
@@
In Fig. 5(m,n), we show a simultaneous study of the skyrmion velocity (Fig. 5(m)) and the magnitude of the variation in radius (Fig. 5(n)), both as a function of the magnitude of the driving microwave magnetic field at a frequency of 57.4 GHz. For details of how the average change in radius was calculated, see supplementary note S4. In Fig. 5(m), at fields above $\sim$20 mT there is a clear linear relationship between the velocity and the driving field. However, below fields of $\sim$20 mT, the skyrmion velocity drops off faster towards 0 m/s, and there is no motion below driving fields of $\sim$3 mT. This region of no skyrmion motion is represented with the grey shaded box. Since this is a simulated system with no pinning, this is clearly an intrinsic effect of their being an insufficient spin-wave emission amplitude to drive motion. Likewise, examining the magnitude of the variation in radius in Fig. 5(n), we draw the same threshold field for skyrmion motion from Fig 5(m) also as a grey shaded box. For these magnetic parameters, we are then able to define a variation in radius of the skyrmions below which there is no motion of $\langle\Delta r_{\mathrm{Sk}}\rangle$ = 0.44 nm, or $\sim$3.5 % of the static skyrmion radius $r_{0}$.
@@
Furthermore, by performing these simulations, we have also developed an understanding of the relationship between skyrmion velocity and microwave power. Figure 5(m) shows the velocity of the skyrmion as a function of the amplitude of the driving field, which is simply a representation of the microwave power driving the spin wave emission. As we have already discussed, for the lowest fields there is no observable skyrmion motion, before a non-linear ramp up in velocity between 3 mT and 20 mT, after which there is a linear relationship between skyrmion velocity and microwave power. This result will help to inform experimental study of skyrmion motion driven by this effect, where the exact relationship between microwave power and magnetic field amplitude will depend on factors such as fabrication of waveguides and losses in connectors, among others.